a boy of weight 25kg slide sdown a rope hanging from the branch of the tree.if the force of friction against him is 50N,the boys acceleration is?

Grade:12

2 Answers

Vikas TU
7 years ago
Simply balance the forces along with friction force.
Therfore,
F – f = ma
250 – 50 = 25a
a = 200/25 =  40/5 = 8 m/s^2
MVRChaitanya
5 years ago
by applying dyanamic equation 
F-fk=ma
mg-50=25a
250-50=25a
200/25=a
so acceleration=8metre/second square.
40/5=a
a=8m/sec2